Chris's wages vary directly as the time she works. If her wages for 20 hours are $150 what is the varation constant? what are the wages for 72.5 hours?
Answer by stanbon(75887)

Chris's wages vary directly as the time she works.
--------
W = kh
-------
If her wages for 20 hours are $150 what is the varation constant?
150= k*20
k = 15/2 = 7.5
--------------
what are the wages for 72.5 hours?
EQUATION:
W = (7.5)(72.5)
W = 543.75
==============
Cheers,
Stan H.